# Quantum Teleportation Coexisting with Conventional Classical Communications in Optical Fiber

### v2

@@
Abstract
The ability for quantum and classical networks to operate in the same optical fibers would aid the deployment of quantum network technology. However, quantum performance can be susceptible to noise photons generated by spontaneous Raman scattering of high-power coexisting classical light. Quantum teleportation is a fundamental operation in quantum networking, but has yet to be demonstrated in fibers populated with high data rate conventional optical signals. In this paper, we demonstrate a three-node quantum state teleportation system coexisting with 400-Gbps C-band classical communications in 30.2 km of fiber. To protect quantum fidelity, Raman noise rates are suppressed using optimized O-band quantum channels and filtering in multiple degrees of freedom. Fidelity is shown to be well maintained with elevated classical powers as high as 18.7 dBm, which could support multiple classical channels with many terabits/s aggregate data rates. These results show the feasibility of advanced quantum and classical network applications operating within a unified fiber infrastructure.
journal: opticajournal
@@ 1 Introduction
Figure 1: (a) Conceptual diagram of the experiment. Alice prepares a qubit $|\psi\rangle_{A}$ that she wishes to transfer onto one photon of an entangled Bell state photon pair at Bob via quantum teleportation. Alice encodes $|\psi\rangle_{A}$ onto a single photon, which is multiplexed into a fiber of length $L_{AC}$ to co-propagate with a high-rate classical communications signal to a node at Charlie. The classical signal by-passes Charlie’s node to continue propagating over a fiber of length $L_{BC}$. Two photons are prepared in an entangled Bell state at Bob’s node. One photon counter-propagates with the classical signal over this fiber to Charlie to undergo a Bell state measurement (BSM) with Alice’s qubit. After a BSM detection, Alice’s state is destroyed whilst the other entangled photon at Bob’s node is projected onto the state $\sigma|\psi\rangle_{A}$, where $\sigma$ is a unitary transformation unique to the BSM result. (b) Experimental implementation. Quantum and classical signals coexist in 30.2  km of fiber ($L_{AC}=15.2\text{ km}$, $L_{BC}=15.0\text{ km}$). Alice and Bob both generate photon pairs with wavelengths of 1290 nm and 1310 nm via non-degenerate spontaneous parametric down-conversion. Alice’s source heralds a single photon, whereas Bob prepares a two-photon entangled state. Qubits are encoded in polarization and the 1290-nm photons from each source are transmitted to Charlie for the BSM. The classical signal has a rate of 400-Gbps and is allocated to the 1547.32 nm C-band wavelength channel. This signal travels an additional 24-km of deployed fiber before and after multiplexing, giving a total classical link distance of 78.2 km (FPC = fiber polarization controller, DWDM = dense wavelength division multiplexer, FWDM = O-band/C-band WDM, PPLN = periodically poled lithium niobate waveguide, $\lambda/2$ = half-wave plate, $\lambda/4$ = quarter-wave plate, LP = linear polarizer, LCR = liquid crystal retarder, PBS = polarizing beam splitter, FBS = 50:50 fiber coupler, VDL = variable optical delay line, FBG = fiber Bragg grating, CIR = circulator, $\text{D}_{j}$ = superconducting nanowire single photon detector, C = common port, P = pass port, R = reflect port, EDFA = erbium doped fiber amplifier).
@@
The optical fiber infrastructure and telecommunications technology that underlie the Internet have been widely adopted by researchers seeking to develop quantum networks capable of applications such as quantum-enhanced cryptography, sensing, and networked quantum computing [1, 2]. However, since the majority of the existing fiber infrastructure is populated with classical network signals, whether or not quantum networking can be realized on a large scale will depend on the ability for both quantum and classical networks to operate in the same fibers due to limited existing fiber connections and the high cost of new installation.
@@
Quantum and classical signals can readily share a single fiber via standard wavelength division multiplexing. However, inelastic scattering of high-power classical light can generate noise photons that obscure the detection of typically sub-photon-level quantum signals. Spontaneous Raman scattering (SpRS) is most prominent due to its broadband spectrum that spans the telecom bands. Without careful design, classical signals require significant attenuation to perform quantum operations, resulting in a trade-off between achievable classical data rates or quantum fidelity.
@@
Quantum-classical integration has been studied for various systems, such as weak coherent states (WCSs) [3, 4, 5, 6, 7, 8, 9, 10, 11, 12], entangled photon pairs [13, 14, 15, 16, 17, 18, 19, 20, 21, 22], Bell state measurements (BSMs) on WCSs [23, 24], continuous variables [25], and squeezed light [26]. However, many systems that are integral to future quantum network architectures, such as those utilizing quantum teleportation, remain unexplored. Using the non-local properties of entanglement, quantum teleportation enables the transfer of a quantum state between two distant physical systems without direct transmission of the systems themselves [27]. It plays a foundational role in advanced applications such as quantum relays [28], quantum repeaters [29], networking quantum computers [30], and other applications in quantum science and technology [31].
@@
Recent years have seen impressive progress towards implementing teleportation-based systems in quantum-dedicated fiber networks [32, 33, 34, 35, 36, 37, 31], but teleportation over fibers carrying conventional classical communications signals has yet to be demonstrated. The ability to perform teleportation alongside modern classical networking without sacrificing fidelity would be an important step in ensuring advanced quantum networks can make full use of the existing fiber infrastructure.
@@
We implement multiple techniques to minimize the degradation of quantum fidelity by SpRS noise. At the quantum source nodes, we generate and distribute photons with O-band wavelengths <1300 nm. Due to the multi-mode SpRS spectrum, these are the telecom wavelengths with nearly the minimum SpRS photon generation probability from the C-band classical source [17]. At quantum receiver nodes, we employ narrow-band spectral filtering and coincidence detection temporal filtering to reject uncorrelated in-band noise photons. This combines various approaches used in high-power quantum-classical coexistence studies of WCSs [9, 10, 11, 12], entangled photons [17], and BSMs on two WCSs [24], but has not been investigated in the context of quantum teleportation.
We first demonstrate that our system’s exceptionally low noise rates allow for no significant impact on entanglement distribution to the BSM nor Hong-Ou-Mandel (HOM) interference [39] between independent photons, which are the underlying operations of the photonic quantum teleportation protocol [38]. We then perform quantum state teleportation, where we measure an average state transfer fidelity of $89.8\pm 3.1$% with elevated classical C-band powers as high as 74 mW (18.7 dBm) launched into the fiber link. We discuss methods to improve both raw teleportation fidelity and noise-resilience as well as the implications of our results for future quantum networking.
@@
This experiment demonstrates approaches for integrating many key classical and quantum network components, including high-rate C-band classical signals and optical amplifiers, single photon and entangled photon pair distribution, and multi-qubit operations such as BSMs all operating simultaneously in the same fibers.
@@ 2 Experimental Design
A conceptual diagram of the experiment is shown in Fig. 1(a). Alice generates a single photon to encode the quantum state $|\psi\rangle_{A}=\alpha|0\rangle+\beta|1\rangle$ that she wishes to teleport to Bob. Alice’s photon is multiplexed into an optical fiber of length $L_{AC}$ to co-propagate with a classical communications signal to a BSM node at Charlie. The classical signal is de-multiplexed just prior to the BSM and then re-multiplexed to bypass Charlie’s node. The classical signal then travels along another fiber of length $L_{BC}$, counter-propagating with respect to one photon from an entangled Bell state photon pair generated at Bob’s node. During the BSM at Charlie, both photons are irreversibly destroyed by the detection whilst Bob’s other photon of the entangled pair is projected onto the state $|\psi\rangle_{B}=\sigma|\psi\rangle_{A}$, where $\sigma$ is a unitary operation that is unique to the BSM result and can be classically communicated to Bob to reconstruct Alice’s state [27].
Fig. 1(b) shows the physical realization of the experiment. In our case, Alice’s state is encoded onto a heralded single photon’s polarization, $L_{AC}$ is 15.2 km, and $L_{BC}$ is 15.0 km. Thus, the quantum state transfer distance, and the distance in which quantum and classical signals coexist, is 30.2 km.
@@
The classical source consists of a 400-Gbps C-band transceiver operating at 1547.32 nm (Ciena WaveLogic 5 Nano 400ZR). To reduce the amount of SpRS noise photons that are generated into the quantum channels, we strategically place our quantum signals in the O-band so that they have an anti-Stokes frequency detuning from the C-band as well as suppressed Raman gain at the far offset detuning between the C- and O-bands [4]. Instead of the commonly used 1310-nm wavelength allocation [9, 10, 11, 12, 24], our transmitted quantum signals are at a wavelength of 1290 nm. This channel has roughly an order of magnitude lower SpRS noise due to the non-uniform SpRS spectrum across the O-band [17], resulting in a higher tolerance to C-band power.
@@
Alice and Bob generate their respective quantum signals via type-0 cascaded second harmonic generation-spontaneous parametric down conversion (c-SHG-SPDC) photon pair generation [40]. Alice and Bob both have independent periodically-poled lithium niobate (PPLN) waveguides that are each phase matched for SHG at 650 nm, which pumps the SPDC process to generate a two-mode squeezed vacuum state with a non-degenerate joint spectrum centered around 1300 nm. This state approximates a single photon pair given the probability of generating a pair per mode is $\ll 1$. The cascaded second order nonlinear process is analogous to four-wave mixing, which we pump using a 1300-nm distributed feedback laser that is intensity modulated using a lithium niobate modulator to generate pulses with an approximate 65-ps temporal full-width at half maximum (FWHM) and a repetition rate of 500 MHz. This pump is then split and directed to Alice and Bob’s nodes.
@@
Alice’s qubit is initially encoded onto a heralded single photon. Her PPLN waveguide is pumped directly to generate photon pairs, where the detection of one photon indicates the presence of a single photon in the other mode. After the waveguide, we use a 100-GHz 1300-nm dense wavelength division multiplexer (DWDM) as a notch filter for the pump, followed by a 1310-nm DWDM to separate each photon into different fibers via the DWDM’s pass and reflect ports. The 1310-nm photon is filtered by a 10.5-GHz FWHM bandwidth fiber Bragg grating (FBG) filter and subsequently detected by a single photon detector. Due to energy conservation, this heralds a single photon with a center wavelength of 1290 nm. Alice’s qubit is then encoded onto the horizontal (H) and vertical (V) components of the heralded photon’s polarization $|\psi\rangle_{A}=\alpha|H\rangle+\beta|V\rangle$ using polarization waveplates. This photon is then multiplexed using a standard O-band/C-band WDM to co-propagate with the 400-Gbps C-band classical signal over 15.2 km of spooled optical fiber (SMF-28(R) ULL) to Charlie.
@@
Bob’s entangled photon pair source is designed similarly, except his PPLN waveguide is placed inside a polarization Sagnac loop to generate the two-photon quantum state $\frac{1}{\sqrt{2}}(|{H_{s}H_{i}}\rangle+e^{i\delta}|{V_{s}V_{i}}\rangle)$. Bob’s 1290-nm photon is transmitted to counter-propagate with the 400-Gbps signal over 15.0 km of spooled fiber (SMF-28(R) ULL) to Charlie. However, Bob’s 1310-nm photon is kept locally to act as the target photon for transferring Alice’s state to Bob’s node. This photon is sent to a free-space polarization module which first compensates rotations in the fibers and prepares the $|\Psi^{-}\rangle_{B}$ Bell state using alignment signals and coincidence detection [41], and then uses a quarter-waveplate, half-waveplate, and polarizing beam splitter (PBS) combination to allow for arbitrary projective measurements of the polarization of Bob’s target photon. Subsequently, this photon is filtered by a 9.8-GHz FBG filter and then detected by a single-photon detector.
Figure 2: (a) Noise photon single photon count rate in the Bell state measurement detectors $D_{1}$ and $D_{2}$ as a function of the 400-Gbps C-band signal’s launch power $P_{\text{cl}}$ injected into the 30.2-km fiber link. (b) Visibility of non-local two-photon interference fringes for Bob’s entanglement distribution over 15.0 km to the Bell state measurement node as a function of $P_{\text{cl}}$ (top) and the measured two-fold coincidence count (CC) interference fringes in the vertical (V) and anti-diagonal (A) bases for $P_{\text{cl}}$ = 74 mW (bottom). (c) Hong-Ou-Mandel interference visibility between heralded photons from Alice and Bob as a function of $P_{\text{cl}}$ (top) and an interference dip measurement when $P_{\text{cl}}$ = 74 mW (bottom).
@@
At Charlie’s node, a BSM is performed on Alice and Bob’s incident polarization qubits using conditional HOM interference at a 50:50 beam splitter [38]. Since HOM interference only occurs between identical photons, the fidelity of the BSM is governed by their spatial and spectro-temporal indistinguishability. To achieve this, a single-mode fiber 50:50 coupler is used for spatial indistinguishability and a variable optical delay in Alice’s pump path is used to ensure that both photons arrive simultaneously at the coupler. Furthermore, the 1290-nm photons are filtered by 10.8-GHz FBGs. Together with the 1310-nm FBGs, this extends the photons’ coherence times relative to the pump’s and increases the probability that each photon is in a single identical spectral mode [42].
These narrow-band spectral filters also play a critical role in improving tolerance to classical power due to the SpRS noise photons being spectrally independent of the quantum signal. Another potential source of noise is residual C-band photons due to insufficient isolation when filtering the O-band quantum signals. This is prevented by placing two cascaded O-band/C-band WDMs before the FBGs and a 1290-nm DWDM, which altogether provides >190 dB isolation from C-band light.
@@
To perform the BSM, we place two PBSs in each output arm of 50:50 coupler before the single photon detectors $D_{1}$ and $D_{2}$, which are set to project onto the orthogonal polarization states $|H\rangle_{D_{1}}$ and $|V\rangle_{D_{2}}$. A two-fold coincidence detection at opposite ports of the 50:50 splitter indicates that Alice and Bob’s photons were in the $|\Psi^{-}\rangle_{AB}$ Bell state due to HOM interference that occurs if they were in any of the three other Bell states [38]. Due to the non-local correlations of Bob’s entangled state $|\Psi^{-}\rangle_{B}$, a successful BSM detection results in the target photon at Bob’s node being equivalent to Alice’s input state $|\psi\rangle_{A}$ [38]. We note that teleportation via detecting $|\Psi^{-}\rangle_{AB}$ can be performed without the PBSs. However, basis-selective Bell state analyzers are most common as they can double the success probability of a BSM by also measuring the $|\Psi^{+}\rangle$ Bell state, which could be done here if detectors were added to the other output ports of the PBSs [43]. Thus, this design allows us to evaluate how coexisting classical signals impact BSM nodes most likely to be implemented in future networks. We also note the PBSs at the BSM naturally filter SpRS noise in the polarization degree of freedom. Due to polarization mode dispersion, SpRS photons generated over long fibers are largely unpolarized [44, 17] which means polarizing elements can reduce noise rates by roughly a factor of two in each detector.
@@
Since the modulation of the classical signal is uncorrelated with the pulsed quantum system, SpRS photons are randomly distributed in time relative to the quantum signals. As a result, narrow detection windows around the expected quantum signal arrival times act as a temporal filter on the noise. Furthermore, the inherent temporal correlations in photon pair detection can provide another layer of timing information that can increase the signal-to-noise ratio compared to sources such as WCSs [45].
The quantum teleportation-classical coexistence study is conducted in a laboratory located at Northwestern University’s Evanston campus. However, the classical transceiver is located in a building at the Northwestern University Chicago campus, which requires transmission over a 24-km deployed fiber pair to and from the Evanston location to operate. Thus, the classical signal travels an additional 48 km of fiber giving a total link distance of 78.2 km. The transmission loss over the 30.2-km fiber is 4.9 dB at the classical signal’s 1547.32-nm wavelength and the 1290-nm quantum signals have a total loss of 10.1 dB (5.1 dB and 5.0 dB for Alice and Bob’s photons, respectively). At Charlie, the bypassing of the BSM by the classical signal using WDMs adds an additional 1.2 dB insertion loss.
The total loss when transmitting the classical signal over the full 78.2-km fiber is 22.8 dB. Due to a receiver sensitivity of -18 dBm, the transceiver’s output power of -9 dBm requires amplification for the system to operate error-free. To achieve this, we place two cascaded erbium doped fiber amplifiers (EDFA) at the Evanston location just before multiplexing into the 30.2-km link to co-propagate with Alice’s photon. The minimum power launched into the 30.2-km fiber to achieve error-free communications could be as low as $-3$ dBm (0.5 mW) for the receiver located 30.2 km + 24 km away from the EDFA, and -10.7 dBm (85 $\mu$W) if the receiver were instead located directly at the end of the 30.2-km link. However, in this study we explore elevated launch powers to emulate more hostile environments that could be encountered in the field such as those found in high-capacity long-distance DWDM classical network systems.

### v3

# Quantum Teleportation Coexisting with Classical Communications in Optical Fiber

@@
Quantum Teleportation Coexisting with Classical Communications in Optical Fiber
[…]
journal: opticajournalarticletype: Research Article
{abstract*}
The ability for quantum and conventional networks to operate in the same optical fibers would aid the deployment of quantum network technology on a large scale. Quantum teleportation is a fundamental operation in quantum networking, but has yet to be demonstrated in fibers populated with high-power conventional optical signals. Here we report to the best of our knowledge the first demonstration of quantum teleportation over fibers carrying conventional telecommunications traffic. Quantum state transfer is achieved over a 30.2-km fiber carrying 400-Gbps C-band classical traffic with a Bell state measurement performed at the fiber’s midpoint. To protect quantum fidelity from spontaneous Raman scattering noise, we use optimal O-band quantum channels, narrow spectro-temporal filtering, and multi-photon coincidence detection. Fidelity is shown to be well maintained with an elevated C-band classical power of 18.7 dBm, which could support multiple classical channels totaling many terabits/s aggregate data rates. These results show the feasibility of advanced quantum and classical network applications operating within a unified fiber infrastructure.
@@ 1 Introduction
The optical fiber infrastructure and telecommunications technology that underlie the Internet have been widely adopted by researchers seeking to develop quantum networks capable of applications such as quantum-enhanced cryptography, sensing, and networked quantum computing [1, 2]. However, since the majority of the existing fiber infrastructure is populated with conventional telecommunications traffic and due to the economic cost of leasing or installing new fiber, whether quantum networking can be realized on a large scale will depend on the ability to propagate quantum signals in the same fiber as high-power classical signals.
@@
Quantum and classical signals can readily share a single fiber via wavelength division multiplexing. However, it was first shown in [3] that noise photons from inelastic scattering of the high-power classical light can obscure the detection of often sub-photon-level quantum signals. Of the potential sources of noise, spontaneous Raman scattering (SpRS) is most dominant due to its broadband spectrum meaning noise can never be entirely prevented [4]. Without careful design, a trade-off between the capabilities of conventional and quantum networking arises.
@@
The study of quantum-classical coexistence has a long history with many experiments conducted for applications using weak coherent state (WCS) sources [3, 4, 5, 6, 7, 8, 9, 10, 11, 12], entangled photon pairs [13, 14, 15, 16, 17, 18, 19, 20, 21, 22], Bell state measurements (BSMs) on WCSs [23, 24], continuous variables [25], and squeezed light [26]. Although significant achievements and insights have been made in these studies, all experiments to date have focused on systems that directly transmit quantum information between network nodes. However, many next-generation quantum applications require the disembodied transfer of quantum states between users. Enabled by the non-local properties of quantum entanglement, quantum teleportation allows the transfer of a quantum state between two distant physical systems without the need for direct transmission [27]. It plays a foundational role in advanced applications such as quantum relays [28], quantum repeaters [29], networking quantum computers [30], and other applications in quantum science and technology [31].
@@
Recent years have seen impressive progress towards implementing teleportation-based networks in quantum-dedicated fiber [32, 33, 34, 35, 36, 37, 38, 39, 31], but teleportation over fibers carrying conventional communications has yet to be demonstrated. As a result, many open questions on the feasibility, limitations, and potential benefits of teleportation in quantum-classical networking remain unexplored.
@@
We implement multiple techniques to mitigate degradation due to noise photons. At the quantum source nodes, we make use of the low SpRS noise in the O-band from strong C-band light [3], which has been shown successful for protecting quantum states from high-power C-band communications [9, 10, 11, 12, 24, 17]. More specifically, we allocate coexisting photons to the 1290-nm quantum channels to have nearly the lowest SpRS generation probability achievable using telecom-band wavelength engineering [17]. At quantum receivers, we use narrow-band spectro-temporal filtering to reject the detection of uncorrelated SpRS photons. Furthermore, each source uses spontaneous parametric down-conversion to generate photon pairs with strong temporal correlations to further reject noise via coincidence detection of the four photons. The design is also similar to a quantum relay, which has been predicted to reduce the impact of detector dark count noise [28] indicating a potential application in quantum-classical coexistence.
We first demonstrate that our approach allows high-power C-band communications to have no significant impact on entanglement distribution to the BSM node nor Hong-Ou-Mandel (HOM) interference [41] between independent heralded photons, which are the underlying operations of the quantum teleportation protocol [40]. We then perform quantum state teleportation and demonstrate that fidelity is maintained at elevated C-band launch powers as high as 74 mW (18.7 dBm), which is on the order of the highest achieved using direct transmission of quantum states [10, 11, 12, 17, 25]. Using a theoretical model, we then discuss the the pros and cons of using noisier quantum wavelength channels (eg., C-band), the effects of storage in imperfect quantum memories, and compare to methods using direct qubit transmission.
@@
This experiment demonstrates approaches for integrating many key classical and quantum network components, including high-rate classical signals, optical amplifiers, single and entangled photon pair distribution, and multi-qubit operations such as BSMs all operating simultaneously in the same fibers.
Figure 1: (a) Conceptual diagram of the experiment. Alice prepares a qubit $|\psi\rangle_{A}$ that she wishes to transfer onto a photon at Bob’s node via quantum state teleportation. Alice encodes $|\psi\rangle_{A}$ onto a single photon, which is wavelength-division multiplexed into a fiber of length $L_{AC}$ to co-propagate with conventional classical communications traffic to a node at Charlie. The classical signal is de-multiplexed to bypass Charlie’s node and re-multiplexed to continue propagating over a fiber of length $L_{BC}$. Two photons generated at Bob’s node are prepared in an entangled Bell state, where one is transmitted to counter-propagate with the classical light to Charlie to undergo a Bell state measurement (BSM) with Alice’s qubit. After a BSM detection, Alice’s state is destroyed whilst the photon kept at Bob’s node is projected onto the state $\sigma|\psi\rangle_{A}$, where $\sigma$ is a unitary Pauli transformation unique to the BSM result. (b) Experimental implementation. Quantum and classical signals coexist in 30.2  km of fiber ($L_{AC}=15.2\text{ km}$, $L_{BC}=15.0\text{ km}$). Alice and Bob both generate photon pairs via non-degenerate spontaneous parametric down-conversion with center wavelengths of 1290 nm and 1310 nm. Alice’s source heralds a single photon, whereas Bob prepares a two-photon entangled Bell state. Qubits are encoded in polarization and the 1290-nm photons from each source are transmitted to Charlie for the BSM and teleportation is evaluated via four-fold coincidence detection and polarization analysis at Bob. The classical signal has a rate of 400-Gbps and is allocated to the 1547.32 nm C-band wavelength channel. This signal travels an additional 24-km of deployed fiber before and after multiplexing giving a total classical link distance of 78.2 km (FPC = fiber polarization controller, DWDM = dense wavelength division multiplexer, FWDM = O-band/C-band WDM, PPLN = periodically poled lithium niobate waveguide, $\lambda/2$ = half-wave plate, $\lambda/4$ = quarter-wave plate, LP = linear polarizer, LCR = liquid crystal retarder, PBS = polarizing beam splitter, FBS = 50:50 fiber coupler, VDL = variable optical delay line, FBG = fiber Bragg grating, CIR = circulator, $\text{D}_{j}$ = superconducting nanowire single photon detector, C = common port, P = pass port, R = reflect port, EDFA = erbium doped fiber amplifier).
@@ 2 Experimental Design
A conceptual diagram of the experiment is shown in Fig. 1(a). Alice generates a single photon to encode the quantum state $|\psi\rangle_{A}=\alpha|0\rangle+\beta|1\rangle$ that she wishes to teleport to Bob. Alice’s photon is multiplexed into an optical fiber of length $L_{AC}$ to co-propagate with a classical communications signal to a BSM node at Charlie. The classical signal is de-multiplexed just prior to the BSM and then re-multiplexed to bypass Charlie’s node. The classical signal then travels along another fiber of length $L_{BC}$, counter-propagating with respect to one photon from an entangled Bell state photon pair generated at Bob’s node. During the BSM at Charlie, both photons are irreversibly destroyed by the detection whilst Bob’s other photon of the entangled pair is projected onto the state $|\psi\rangle_{B}=\sigma|\psi\rangle_{A}$, where $\sigma$ is a unitary operation that is unique to the BSM result and can be classically communicated to Bob to recover Alice’s state [27] either by physically applying the unitary or accounting for it in data post-processing. Fig. 1(b) shows the physical realization of the experiment. In our case, Alice’s state is encoded onto a heralded single photon’s polarization, $L_{AC}$ is 15.2 km, and $L_{BC}$ is 15.0 km. Thus, the quantum state transfer distance, and the distance in which quantum and classical signals coexist, is 30.2 km.
@@
The classical source consists of a 400-Gbps C-band transceiver operating at 1547.32 nm (Ciena WaveLogic 5 Nano 400ZR). To reduce the amount of SpRS noise photons that are generated into the quantum channels, we generate photons at O-band wavelengths so that they have an anti-Stokes frequency detuning from the C-band light as well as suppressed Raman gain at the far offset detuning between the C- and O-bands [4]. Instead of the most commonly used 1310-nm O-band channel [9, 10, 11, 12, 24], we suppress SpRS by roughly another order of magnitude using the 1290-nm channel due to the multi-mode SpRS noise spectrum from C-band light [17].
@@
Both Alice and Bob generate their respective quantum signals via type-0 cascaded second harmonic generation-spontaneous parametric down conversion (c-SHG-SPDC) photon pair generation [42]. Alice and Bob each have independent periodically-poled lithium niobate (PPLN) waveguides that are phase matched for SHG at 650 nm. The cascaded second-order nonlinear process is analogous to four-wave mixing, which we pump using a 1300-nm continuous-wave distributed feedback laser that is intensity modulated using a lithium niobate Mach-Zehnder modulator to generate pulses with an approximate 65-ps temporal full-width at half maximum (FWHM) and a pulse repetition rate of 500 MHz. This pump is then split and directed to Alice and Bob’s nodes (see supp. materials). The 650-nm pulses generated inside each waveguide simultaneously drives SPDC to produce a two-mode squeezed vacuum state with a non-degenerate joint spectrum centered around 1300 nm. This state approximates a probabilistic source of single-photon pairs given the probability of generating a pair per pulse is $\ll 1$.
@@
Alice’s qubit is initially encoded onto a heralded single photon. Her PPLN waveguide is pumped directly to generate photon pairs. After the waveguide, we use a 100-GHz 1300-nm dense wavelength division multiplexer (DWDM) as a notch filter for the pump, followed by a 1310-nm DWDM to separate each photon into different fibers via the DWDM’s pass and reflect ports. The 1310-nm photon is filtered by a fiber Bragg grating (FBG) with FWHM bandwidth of 60 pm and then detected by a superconducting nanowire single-photon detector (SNSPD). Due to energy conservation, this detection heralds the presence of a photon with a center wavelength of 1290 nm in the other mode. Alice’s qubit is then encoded onto the horizontal (H) and vertical (V) components of the heralded photon’s polarization $|\psi\rangle_{A}=\alpha|H\rangle+\beta|V\rangle$ using polarization waveplates. This photon is then multiplexed using an O-band/C-band WDM to co-propagate with the 400-Gbps C-band classical signal over 15.2 km of spooled optical fiber (SMF-28(R) ULL) to Charlie.
@@
Bob’s source is designed similarly, except his PPLN waveguide is placed inside a polarization Sagnac loop to generate polarization entangled photon pairs using c-SHG-SPDC [42]. Bob’s 1290-nm photon is transmitted to counter-propagate with the 400-Gbps signal over 15.0 km of spooled fiber (SMF-28(R) ULL) to Charlie. However, Bob’s 1310-nm photon is kept locally to act as the target photon for teleporting Alice’s state to Bob’s node. This photon is sent to a free-space polarization module which compensates polarization rotations in the fibers and prepares Bob’s photons in the $|\Psi^{-}\rangle_{B}$ Bell state using alignment signals and coincidence detection [43], and then uses a quarter-waveplate, half-waveplate, and polarizing beam splitter (PBS) combination to characterize the polarization of Bob’s target photon. Subsequently, this photon is filtered by a 56 pm bandwidth FBG filter and then detected by an SNSPD.
@@
At Charlie’s node, a BSM is performed on Alice and Bob’s incident polarization qubits using conditional HOM interference at a 50:50 beam splitter [40]. Since HOM interference only occurs between identical photons, the fidelity of the BSM is governed by their spatial and spectro-temporal indistinguishability. To achieve this, a single-mode fiber 50:50 coupler is used for spatial indistinguishability and a variable optical delay in Alice’s pump path is used to ensure that both photons arrive simultaneously at the coupler. Furthermore, the 1290-nm photons are filtered by 60-pm bandwidth FBGs. Together with the 1310-nm FBGs, this extends the photons’ coherence times relative to the pump’s to increase the probability that each photon is in a single identical spectral mode [44]. These narrow-band filters are also crucial for improving tolerance to classical power due to the spectral independence between the SpRS noise photons and the quantum signals. To prevent the detection of C-band photons due to insufficient filter isolation, we cascade two cascaded O-band/C-band WDMs before the FBGs and a 1290-nm DWDM to achieve >190 dB rejection of C-band light.
@@
To perform the BSM, we place PBSs in each output of the 50:50 coupler before the detectors $D_{1}$ and $D_{2}$ which are set to project onto orthogonal polarizations $|H\rangle_{D_{1}}$ and $|V\rangle_{D_{2}}$. A two-fold coincidence detection then indicates that Alice and Bob’s photons were in the $|\Psi^{-}\rangle_{AB}$ Bell state due to HOM interference effects on the other Bell states [40]. Due to the non-local correlations of Bob’s entangled photons in the state $|\Psi^{-}\rangle_{B}$, the destructive measurement of $|\Psi^{-}\rangle_{AB}$ results in the remaining photon at Bob’s node being projected onto Alice’s initial prepared state $|\psi\rangle_{A}$ [40]. Although teleportation can be performed without projecting onto the $H/V$ basis, Bell state analyzers that measure in the qubit’s eigenbasis can double the success probability of a BSM by also measuring the $|\Psi^{+}\rangle$ Bell state if detectors were added to the other ports of our PBSs [45]. Including the PBSs thus allows us to investigate BSM nodes most likely to be implemented in future networks. We also note that due to polarization mode dispersion over long fibers, the SpRS noise is largely unpolarized [46, 17] which means the polarizing elements roughly halve the amount of SpRS entering each detector.
@@
Since the classical data modulation is uncorrelated with the pulsed quantum sources in the time-domain, the coincidence detection window acts as a temporal filter of the SpRS noise. The temporal correlations in photon pair generation can also be exploited for time-domain filtering. Heralding single photons can improve tolerance to noise compared to sources such as WCSs by accepting a detection only when the heralding signal ensures a photon had been generated which reduces the probability of transmitting a vacuum state [47]. In our four-photon teleportation system, both Alice and Bob’s pair sources benefit from these correlations and midpoint BSM nodes in a quantum relay configuration have been predicted to offer higher tolerance to dark count effects [28]. Given this, we predict that instead of Alice directly transmitting her qubit to Bob, our system can potentially offer higher tolerance to classical power in quantum-classical networks (see supp. materials).
The coexistence study is conducted over spooled fiber in a laboratory located at Northwestern University’s Evanston campus. The transmission loss over the 30.2-km fiber is 4.9 dB at the classical signal’s 1547.32-nm wavelength and the 1290-nm quantum signals have a total loss of 10.1 dB (5.1 dB and 5.0 dB for Alice and Bob’s photons, respectively). At Charlie, the bypassing of the BSM by the classical signal using WDMs adds an additional 1.2 dB insertion loss. However, the classical transceiver is located at Northwestern’s Chicago campus which is connected to Evanston by a 24 km deployed fiber pair and thus operates over a total link distance of 30.2 + 48 km = 78.2 km. Due to the high loss over the full link (22.8 dB) and an initial output power of -9 dBm, amplification is needed to meet the classical system’s $-18$ dBm minimum received power requirements. We achieve this using erbium doped fiber amplifiers (EDFAs) placed just before multiplexing to co-propagate with Alice’s photon. The minimum power launched into the 30.2-km fiber to operate the 400-Gbps channel is $P_{\text{min}}=-3$ dBm (0.5 mW) for the receiver located 30.2 km + 24 km away from the EDFA, and -10.7 dBm (85 $\mu$W) if the classical system operated solely over the 30.2-km link. However, we explore elevated powers using the EDFAs to evaluate whether teleportation can coexist with more hostile classical systems such as ultra high-power long-distance DWDM conventional communications.
Figure 2: (a) SpRS noise photon count rate in the Bell state measurement (BSM) detectors $D_{1}$ and $D_{2}$ as a function of the classical launch power $P_{\text{cl}}$ and the ratio $M$ to the minimum power $P_{\text{min}}$ needed to operate a single 400-Gbps channel. (b) Bob’s entanglement distribution visibility $V_{\rm ent}$ over 15.0 km to the BSM node versus $P_{\text{cl}}$ (top) and the corresponding two-fold coincidence count (CC) non-local interference fringe measured in the vertical (red) and anti-diagonal (blue) bases for $P_{\text{cl}}$ = 74 mW (bottom). (c) Hong-Ou-Mandel interference visibility $V_{\rm HOM}$ over 30.2 km between heralded photons from Alice and Bob’s sources versus $P_{\text{cl}}$ (top) and the corresponding four-fold coincidence interference fringe when $P_{\text{cl}}$ = 74 mW (bottom). The dashed lines show the predictions of our theoretical model [48] (see supp. materials).